Avicorp has a $12.1 million debt issue​ outstanding, with a 5.8% coupon rate. The debt has​ semi-annual coupons, the next coupon is due in six​ months, and the debt matures in five years. It is currently priced at 94% of par value.

a. What is​ Avicorp's pre-tax cost of​ debt? Note: Compute the effective annual return.

b. If Avicorp faces a 40% tax​ rate, what is its​ after-tax cost of​ debt?

​Note: Assume that the firm will always be able to utilize its full interest tax shield.

a. The cost of debt is

nothing​%

per year.  ​(Round to four decimal​ places.)b. If Avicorp faces a40%

tax​ rate, the​ after-tax cost of debt is

nothing​%.

​(Round to four decimal​ places.)
Answer #1

a. Price =94%*1000 =940
Semi annual coupon =5.8%*1000/2 =29
Number of Periods =5*2 =10
Par Value =1000
Pre tax semiannual cost of Debt =RATE(10,29,-940,1000) =3.62604%
Pre tax effective rate =(1+3.6204%)^2-1 =0.073719 or 0.0737

b. After tax cost of debt =0.073719*(1-40%) =5.1603% or 0.0516